OpenStudy (anonymous):

find the required value by setting up the general equation and then evaluating. find r when n=160 if r varies directly as the square root of n and r=4 when n=250

OpenStudy (anonymous):

igo t .25 and then .25(160=40 is this correct?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

t is the letter you're using for your constant of variation?

zepdrix (zepdrix):

`r varies directly as the square root of n`\[\Large r\quad=\quad t\;\sqrt{n}\]

zepdrix (zepdrix):

Plugging in the information they gave us:\[\Large 4\quad=\quad t\;\sqrt{250}\]To t works out to be uhhhhhhh,\[\Large t\quad=\quad 0.25298\] Ah good looks like you're on the right track! :)

zepdrix (zepdrix):

\[\Large r \quad=\quad .25\;\sqrt{n}\] `find r when n=160` \[\Large r\quad=\quad .25\;\sqrt{160}\]

zepdrix (zepdrix):

Your last step looks a little bit off. Hmm.

OpenStudy (anonymous):

r=3.2

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

lol thank you